Hawaii-2016-08 experiment with new MOBYrefresh blue Resonon - BS04cfg001

First data set with the BS04 which will eventually go into MOBY262. Mike tooks some lamp data which can be used to check from track movement with data taken at Resonon before shipment. Also Mike is starting the fiber splitter temperature tests Ken requested. For the day 2 data the Ambient %RH/degC pen was inside the oven next to the splitter, and the External thermistor was also inside the oven, screwed to the side of the splitter. Day 2 the spliiter died at the end of the testing. Not sure why.

Problems seen (and solved) and good things:

  1. Splitter died - day02 PageNum 3.01. After turning the oven on for a few minutes and taking data which all went fine. Mike opened the oven door to cool off the oven and the Solis software froze and at the same time the splitter died and the lite scan look like darks.
  2. Splitter response to a temperature change from ~75F to 110F for 13 files has a percent standard deviation of less than 0.2%

Log Sheets

Photos

All the raw data files and their images , Table of KEYWORDS variable
Page Number
Link
Description
Date
1.01
Day 01 - 13 Sep Raw Data , Table
DAY 1 BS04 - Oriel, fiber splitter test
Sep 13, 2016
1.02
Day 02 - 14 Sep Raw Data , Table
DAY 2 BS04 - Oriel, fiber splitter test, in the oven (jumper fibers added to output fibers)
Sep 13, 2016
 
Day01 - First data set with BS04 - Oriel, fiber splitter test
2.01 Track definitions Checking Track definitions from Resonon data Sep 13, 2016
2.02 Track movement Looking at track movement from Resonon -> Hawaii Sep 13, 2016
2.03 Lamp data Looking a the Track 7 lamp data Sep 13, 2016
2.04 Splitters First look at fiber spliter test data Sep 14, 2016
DAY 2 BS04 - Oriel, fiber splitter test, in the oven
3.01 Spliiter w/ temp change First look at fiber splitters in the oven Sep 14, 2016
       
       

 

DAY 1 EMAIL from Mike

Hi Steph,
I'm FTPing you some MORE BS DATA! I just realized I am putting the .fits files in the wrong place - I did not make a "day01" subdir under /ftp1/Mike/HI-2016-08/BS04data/ Sorry!  Could you straighten that out for me please?

These would be our first-data via BS04cfg001, which should be the instrument to deploy on MOBY262.

I began with 3x repeats of bac/sig via the Oriel incandescent + sphere at Andor TEC=off, -20, -40, -80 degC w/ all bare bundle FOs @ sphere. This was to look for bad pixels (I didn't see any, and, I am still working on that 16346 hot pixel camera data, so I am behind on that now...).

Some of these should also serve as a track-edge-selection data...

Next I set the TEC @ -40 degC, and coupled a new FO splitter to bundle FO's 5 & 10, plus I attached the spare external thermistor to the splitter - I think this is on AD chan #5 if the 5 Vdc source is counted as chan #0 (?).  It looks like what I called splitter output fiber #B @ track 10 was the brighter FO, so maybe I will re-name this splitter fiber #1 - so #1 = brighter & #2 = dimmer (?).

There was an hour delay between these 2 sub-sets of data where I numbered the bundle FOs and I dug out some jumper FOs from the double-stacked L-vans out back (ho, is that one BEEEG MESS!).

I have high hopes to put this splitter in the oven tomorrow and collect some thermal data for Kens' report (that's where we'll want to know which is the AD external thermistor, and how to get Celsius degrees out of it - I guess I can also put the %RH/degC probe in the oven too...).

Thanks, and good luck, MF

DAY 2 EMAIL from Mike

Hi Steph,
Well, there are 14x bac/sig .fits pairs heading your way, plus 2x log sheet scans and 1x jumper FO spec sheet scan, plus 1x ad data .dat and 1x pd data .txt, ( I changed the ad file name from "aux*" to "ad*" to match "pd*"...) plus 10x photos of the setup.

This was the first oven-baked fiber-splitter experiment. I used the first Serial Number of the new splitter batch.

However, I think I killed the splitter.  It no longer passes light... I'm rather confused aboot this outcome, because the Solis software kinda froze up, such that I had to restart it, and reset the camera TEC to -40 deg C, justt after I opened and closed the oven door to let in some cool air, then after I restarted Solis & reset the TEC to -40 and waited for that to stabilize there was only dark ADU level via signal scan. I don't think the open/close oven door could have pinched the splitter fibers in any way.  Maybe the "blast" of "cold" air was too traumatic for the splitter?  But why would the Solis acquisition software also freeze at that moment? I have seen Solis freeze up before, but I'm not sure exactly what were the symptoms...

I think I had your oven on, at the minimum dial setting just to get it heating, for a total of two minutes!

I'm curious what the AD temperatures time series looks like! I saw no significant change in signal ADU, but I'm curious if you do. I'm also curious if the splitter comes back to life tonight!
MF